OpenStudy (anonymous):

The sum of the digits of a two-digit number is 12. A new number is formed when the digits are reversed that is 12 less than twice the original number. Find the new number. Can someone lead me in the right direction? I want to find it myself, I'm just not sure of how to do that.. Thanks! :)

OpenStudy (btaylor):

original number: 10x+y (x is tens digit, y is ones digit) new number: 10y+x

OpenStudy (btaylor):

x+y=12 10y + x = 2(10x+y) - 12 it's a system.

OpenStudy (anonymous):

So I just use any two numbers I want? Like 3 & 9?

OpenStudy (btaylor):

no, simplify the second equation into: 10y+x = 20x+2y+12 then solve for x or y (I'll solve for y) 8y=19x+12 no, with the first equation, y = 12-x. So wherever you see a y, replace it with 12-x 8(12-x)=18x+12 so now you can solve for x

OpenStudy (anonymous):

When I solve for x, will that give me one of the numbers? & Thank you for helping :)

OpenStudy (btaylor):

it will give you the x value. then use y=12-x to solve for y. you want to find the new number, so do 10y+x

OpenStudy (anonymous):

Oh okay. Thank you!!